rssLink RSS for all categories
 
icon_orange
icon_green
icon_green
icon_orange
icon_red
icon_green
icon_green
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_red
icon_orange
icon_red
icon_blue
icon_green
icon_orange
icon_orange
icon_green
icon_orange
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_green
icon_green
icon_green
 

FS#37330 — Pop Ashburn

Attached to Project— Reseau Internet et Baies
Incident
Backbone US
CLOSED
100%
We are experiencing a loss of signal on our Dark fiber between Baltimore and Washington.

We have contacted our provider and ask for them to investigate and correct the situation.

Date:  Thursday, 07 March 2019, 23:19PM
Reason for closing:  Done
Additional comments about closing:  Work is finished
Comment by OVH - Wednesday, 06 March 2019, 21:23PM

we isolated Cogent, Equinix and Telefonica to avoid any saturation.


Comment by OVH - Wednesday, 06 March 2019, 21:50PM

Moreover we isolated all the links between Ashburn and Newark, the traffic to Ashburn is currently going through Chicago


Comment by OVH - Thursday, 07 March 2019, 02:22AM

Our Provider OSP tech are doing OTDR traces to pin-point the location.


Comment by OVH - Thursday, 07 March 2019, 03:44AM

Our Provider OSP tech are continuing to perform OTDR testing in an effort to locate damage on this cable. As a temporary fix they are looking into a possible reroute to a different cable.


Comment by OVH - Thursday, 07 March 2019, 06:08AM

Current Update: splicing techs are still in the area continuing to work. They are not hands off yet.


Comment by OVH - Thursday, 07 March 2019, 06:49AM

Fibers between Ashburn and Newark are repaired.
Technicians are still working on Ashburn Beauharnois


Comment by OVH - Thursday, 07 March 2019, 09:35AM

crews remain onsite at this time working to get all services roll to new fiber to restore services


Comment by OVH - Thursday, 07 March 2019, 14:44PM

OSP techs are finishing up the final migrations now. All remaining circuits will have to be repaired for services to restore. Due to the large impact of the cable it may take longer that 07:30 CST to process the emergency maintenance . We will advise as soon as the maintenance has been complete. More updates to follow.


Comment by OVH - Thursday, 07 March 2019, 16:39PM

The emergency maintenance has been complete.
OSP crews will now cut the cable and being working on repairs at that time.


Comment by OVH - Thursday, 07 March 2019, 22:08PM

Our fiber is now repaired and services are restored.

We are waiting confirmation that the work is finished.